OpenStudy (anonymous):

Which of these is not a geometric sequence? 3, 6, 12, 24, ... 250, 50, 10, 2, .... 5, 15, 45, 135, ... 4, 12, 20, 28, ...

OpenStudy (the_fizicx99):

D

OpenStudy (the_fizicx99):

They're increasing by +8

OpenStudy (anonymous):

in the geometric sequence the common ratio ( ratio between two successive terms should be constant)

OpenStudy (the_fizicx99):

Geometric is a ratio, use the first option for example, 24 / 12 = 2 12 / 6 = 2 6 / 3 = 2